# BiCo-Fusion: Bidirectional Complementary LiDAR-Camera Fusion for Semantic- and Spatial-Aware 3D Object Detection

### v1

@@ IV EXPERIMENTS
TABLE I: Performance Comparison of Different Methods. ‘L’ and ‘C’ denote the LiDAR and the camera. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ refer to the construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. ‘†’ denotes the model with test-time augmentation and model ensemble techniques.
@@
BiCo-Fusion (Ours) L+C 76.1 77.1 89.6 67.4 44.1 76.4 66.8 83.5 82.6 65.9 93.6 90.8
Figure 5: Qualitative results of BiCo-Fusion on nuScenes validation set.
@@ IV-A Experimental Setup
Dataset. Following previous works [1, 2], we evaluate the 3D object detection performance of our BiCo-Fusion on the nuScenes benchmark [35]. The nuScenes dataset is a large-scale autonomous driving benchmark including 10,000 driving scenarios in total (700, 150, and 150 scenes for training, validation, and testing, respectively). To evaluate 3D object detection, there are two main metrics: mean Average Precision (mAP) and nuScenes detection scores (NDS). The mAP is calculated from the mean of the average precision across ten classes under distance thresholds of 0.5m, 1m, 2m, and 4m. NDS is a comprehensive metric that integrates mAP with five True Positive (TP) metrics: mATE, mASE, mAOE, mAVE, and mAAE, which together ensure a robust evaluation of object translation, scale, orientation, velocity, and attributes.
@@
TABLE II: Ablation studies for our proposed components on the nuScenes validation set.
@@
TABLE III: Parameter selection in Voxel Enhancement Module.
@@
IV-B Main Results
We compare our BiCo-Fusion with leading methods on the nuScenes test set, as shown in Tab. I. It demonstrates that BiCo-Fusion significantly improves the LiDAR-only baseline, TransFusion-L [2], by 6.9% mAP and 4.3% NDS, due to the additional fusion of multi-modalities. More importantly, BiCo-Fusion achieves superior performance compared to prior works, with 72.4% mAP and 74.5% NDS. By combining test-time augmentation and model ensemble techniques, BiCo-Fusion further achieves a state-of-the-art (SOTA) performance. In Fig. 5, we also present some qualitative results on the nuScenes validation set, which shows that our method can detect challenging cases even when the small objects are occluded or many objects are tightly packed together.
IV-C Ablation Studies
IV-C1 Componet-wise Ablation
As shown in Tab. II, we conduct ablation studies on our proposed modules on the nuScenes validation set. In Tab. II (1), we reimplement TransFusion-L [2] as our LiDAR-only baseline. According to Tab. II (2), after combing VEM, we improve mAP by 1.9% and NDS by 1%, which indicates the effectiveness of enhancing the semantic capability of voxel features. Furthermore, incorporating the Unified Fusion (U-Fusion) leads to additional improvement by 3.5% mAP and 2.2% NDS (Tab. II (3)). It should be emphasized that there are two main reasons why the combination with Unified Fusion contributes to such a big improvement: a) The Voxel Enhancement Module in Pre-Fusion reduces the modality gap, setting the stage for the second step of fusion. b) The fusion in a unified voxel space fills in many non-empty voxels, compensating for the geometric sparsity of the LiDAR modality.
Finally, with all these components combined (Tab. II (4)), the final mAP and NDS achieve significant enhancement to 70.5% and 72.9%, respectively, which demonstrates that enhancing the spatial capability of camera features can further decrease the negative effect and help to guarantee the strong representations can be preserved.
IV-C2 Discussion on Voxel Enhancement Module.
We analysis the selection of the number of surrounding camera features per voxel ($K$) in the voxel enhancement module. As shown in Tab. III, when $K$ increases from 1 to 9, the mAP and NDS gradually rise. And when $K$ is increased to 10, there is no change in NDS, but the mAP decreases, so we select $K=9$ for the framework.
Tab. IV demonstrates the effectiveness of our distance-prior weighting scheme in the voxel enhancement module, which improves the mAP and NDS by 0.8% and 0.6%, respectively.
IV-C3 Discussion on Unified Fusion
In Unified Fusion (U-Fusion), we employ an adaptive weighted fusion method, which dynamically selects features from both modalities for fusing in unified voxel space. We ablate it in Tab. V, which indicates that our scheme improves performance compared with simple concatenation, by 1.1% mAP and 0.7% NDS.
TABLE IV: Effect of Distance-prior weighting scheme in VEM.
@@ IV-C3 Discussion on Unified Fusion
TABLE V: Effect of adaptive weighting in U-Fusion
@@
TABLE VI: Ablation studies for different choices in framework
@@ IV-C4 Ablation on Different Choices in Framework
We further construct ablation studies on different choices in our BiCo-Fusion framework. As shown in Tab. VI (a), our framework can benefit from more powerful image backbones. Swin-T [37] achieves the best performance, with improving mAP by 1.1%, compared with ResNet [52]. Tab. VI (b) demonstrates that a smaller voxel size can bring a small-scale improvement, with about 0.5% for both mAP and NDS. Same as voxel size, using a larger image resolution (e.g., 384 × 1056) also leads to a slight performance improvement, as shown in Tab. VI (c).
V CONCLUSIONS
This letter presented BiCo-Fusion, a novel framework that employs a bidirectional complementary LiDAR-camera fusion strategy to preserve the spatial structure of LiDAR and the semantic features of the camera, ensuring a reduction of the modality gap for robust 3D object detection. We introduced the Pre-Fusion and Unified Fusion (U-Fusion) to enhance the semantics of LiDAR voxel features and spatial capability of camera features and adaptively fuse the enhanced features from both modalities in a unified voxel space. Extensive experiments demonstrated the effectiveness of these components and our proposed framework.

### v3

@@ IV EXPERIMENTS
TABLE I: Comparison with SOTA methods on nuScenes test set. ‘L’ and ‘C’ denote the LiDAR and the camera. ‘C.V.’, ‘Motor.’, ‘Ped.’, and ‘T.C.’ refer to the construction vehicle, motorcycle, pedestrian, and traffic cone, respectively. ‘†’ denotes the model with test-time augmentation and model ensemble.
@@
BiCo-Fusion† (Ours) L+C 76.1 77.1 89.6 67.4 44.1 76.4 66.8 83.5 82.6 65.9 93.6 90.8
Figure 5: Comparison of the qualitative results between our method and others on nuScenes validation set. We provide results of ground truth, TransFusion, and our BiCo-Fusion. Results on cars are colored in yellow, pedestrians in blue, and trucks in red. The red circles highlight the difference among these results. We can observe BiCo-Fusion can detect more occluded objects and some small objects, which indicates the superiority of our method.
@@ IV-A Experimental Setup
Dataset. Following[1, 2], we evaluate the 3D object detection performance of our BiCo-Fusion on the nuScenes benchmark [36]. The nuScenes dataset is a large-scale autonomous driving benchmark including 1,000 driving scenarios in total (700, 150, and 150 scenes for training, validation, and testing, respectively). To evaluate 3D object detection, there are two most important metrics: mean Average Precision (mAP) and nuScenes detection scores (NDS), which reflect the total performance of a method. The mAP is calculated from the mean of the average precision (AP) across ten classes under distance thresholds of 0.5m, 1m, 2m, and 4m. NDS is a robust metric that integrates mAP with five True Positive (TP) metrics: mATE, mASE, mAOE, mAVE, and mAAE, which together ensure a comprehensive evaluation of object translation, scale, orientation, velocity, and attributes.
TABLE II: Comparison on the nuScenes validation set.
     Method      mAP      NDS      FPS
     BEVFusion [35]      67.9      71.0      0.8
     TransFusion [2]      67.5      71.3      3.4
     BEVFusion [1]      68.5      71.4      4.3
     BiCo-Fusion (Ours)      70.5      72.9      2.8
@@
TABLE III: Ablation studies for our proposed components on the nuScenes validation set.
@@
We additionally combine BiCo-Fusion with test-time augmentation (TTA) and model ensemble techniques [6] to get the BiCo-Fusion†. Specifically, our BiCo-Fusion† ensembles several BiCo-Fusion models with different voxel sizes ranging from (0.05m, 0.05m, 0.2m) to (0.125m, 0.125m, 0.2m), with intervals of 0.025m. Each model is adopted with standard test-time augmentation like double flipping and rotations (i.e., 0°, ±22.5°, ±180°) to the input point clouds. About the details on how to combine TTA and model ensemble to produce the final detection, our implementation is consistent with prior works [6, 47, 1, 38]. Specifically, we first train these multiple models with different voxel sizes. During inference, TTA is applied to each model, where the input data undergoes the above transformations to generate multiple augmented samples. Then each model produces predictions for these augmented samples. Following [38, 6], we use weighted box fusion (WBF) [54] to fuse these predictions, obtaining a single prediction per model. Finally, for model ensemble, we use non-maximum suppression (NMS) [55] to ensemble predictions from all models to produce the final result.
IV-B Main Results
In the fair setting without using TTA and model ensemble, we compare our BiCo-Fusion with leading methods on nuScenes test set, as shown in Tab. I. The class-specific APs of BiCo-Fusion stay very close to some very recent works, including [38] and [27]. More importantly, BiCo-Fusion improves the LiDAR-only baseline [2] by 6.9% mAP and 4.3% NDS with the additional fusion of multi-modalities, achieving superior performance with 72.4% mAP and 74.5% NDS, which demonstrates we outperform all existing SOTA 3D detection methods with better average performance across all classes, and better overall performance under comprehensive evaluation setting of the NDS [36]. As an example, although CMT [38] gets the highest APs in some classes (Truck, Bus, Bike), our method outperforms it in more classes (Car, C.V., Trailer, Barrier, Ped., T.C.), which makes our method superior in average performance. By simply adding TTA and model ensemble techniques, which are widely adopted by prior works for submissions on the nuScenes leaderboard [6, 1, 40, 47, 34, 38], the BiCo-Fusion† version achieves a huge performance improvement with 76.1% mAP and 77.1% NDS, reflecting the scalability of our method.
TABLE IV: Parameter selection in Voxel Enhancement Module.
@@ IV-B Main Results
TABLE V: Effect of Distance-prior weighting scheme in VEM.
@@
As shown in Tab. II, BiCo-Fusion achieves superior 3D detection performance on the nuScenes validation set while maintaining a comparable inference speed. Fig. 5 presents the qualitative results on the nuScenes validation set. Our method predicts more accurate results in some complex cases like the objects are occluded or many objects are tightly packed together.
IV-C Ablation Studies
IV-C1 Component-wise Ablation
As shown in Tab. III, we conduct ablation studies on our proposed modules on the nuScenes validation set. In Tab. III (1), we reimplement TransFusion-L [2] as our LiDAR-only baseline. According to Tab. III (2), after combining VEM, we improve mAP by 1.9% and NDS by 1%, which indicates the effectiveness of enhancing the semantic capability of voxel features. In Tab. III (3), combining IEM and U-Fusion, the mAP and NDS improve to 69.2% and 72.0%, respectively, which brings a large shift. But the overall performance is lower than the combination of VEM and U-Fusion, suggesting that enhancing the semantic awareness of the LiDAR modality (i.e., VEM) leads to better improvement than enhancing the 3D spatial awareness of the camera modality (i.e., IEM). Incorporating the Unified Fusion (U-Fusion) with VEM (Tab. III (4)) leads to additional improvement by 3.5% mAP and 2.2% NDS compared to using only VEM. It should be emphasized that there are two main reasons why the combination contributes to such a big improvement: a) The Voxel Enhancement Module in Pre-Fusion prepares more comprehensive LiDAR features, setting the stage for the second step of fusion. b) The fusion in a unified voxel space fills in many empty voxels, compensating for the geometric sparsity of the LiDAR modality.
Finally, with all these components combined (Tab. III (5)), the final mAP and NDS achieve significant enhancement to 70.5% and 72.9%, respectively, demonstrating the effectiveness of our whole framework, which ensures semantic- and spatial-aware 3D object detection.
IV-C2 Discussion on Voxel Enhancement Module
We analyze the selection of the number of surrounding camera features per voxel ($K$) in the voxel enhancement module. As shown in Tab. IV, when $K$ increases from 1 to 9, the mAP and NDS gradually rise. And when $K$ is increased to 10, there is no change in NDS, but the mAP decreases, so we select $K=9$ for the framework.
Tab. V demonstrates the effectiveness of our distance-prior weighting scheme in the voxel enhancement module, which improves the mAP and NDS by 0.8% and 0.6%, respectively.
IV-C3 Discussion on Unified Fusion
In Unified Fusion (U-Fusion), we employ an adaptive weighting scheme, which dynamically fuses features from both modalities to build a unified representation. We ablate it in Tab. VI, which indicates that our scheme improves performance compared with simple concatenation, by 1.1% mAP and 0.7% NDS.
TABLE VI: Effect of adaptive weighting in U-Fusion.
@@ IV-C3 Discussion on Unified Fusion
TABLE VII: Ablation results under different settings.
@@ IV-C4 Ablation on Different Choices in Framework
We further construct ablation studies on different choices in our BiCo-Fusion framework. As shown in Tab. VII (a), our framework can benefit from more powerful image backbones. Swin-T [41] achieves the best performance, with improving mAP by 1.1%, compared with ResNet [56]. Tab. VII (b) demonstrates that a smaller voxel size can bring a small-scale improvement, with about 0.5% for both mAP and NDS. Same as voxel size, using a larger image resolution (e.g., 384 × 1056) also leads to a slight performance improvement, as shown in Tab. VII (c).
V CONCLUSION
This letter presented BiCo-Fusion, a novel LiDAR-camera fusion framework for achieving semantic- and spatial-aware 3D object detection. In BiCo-Fusion, we introduced Pre-Fusion with a bidirectional complementary strategy to enhance the semantic awareness of LiDAR features and the 3D spatial awareness of camera features, which ensures comprehensiveness at the semantic and spatial levels for both modalities, and also Unified Fusion (U-Fusion) to adaptively fuse the enhanced features from both modalities for building a unified representation. Extensive experiments demonstrated the effectiveness of these components and our proposed framework.